Reports 1-1 of 1 Clear search Modify search
AdV-DET (Commissioning)
bonnand, tournefier. - 19:07 Thursday 03 June 2021 (52001) Print this report
SDB2 adaptation to high power and beam dump installation.

Here is a summary of the work on SDB2 from tuesday morning the 1st to thursday afternoon the 3rd.

 

    Weight table
    Elements Removed (g) Added (g)
    Door old 6144 + 6161 = 12305  
    New doors   3830 + 3854 = 7684
    new bridge   777
    old bridge (+connector and screw) 533  
    Air box 25 (old B5_PD2) 854  
    Air box 56 (new B5_PD2)   901
    Air box 01 (old B1p_PD2) 857  
    Air box 63 (new B1p_PD2)   882
    Beam dump on PD (11 g each) x 6   66
    1 inch mount B1p_M1b 157  
    1 inch mount B1p_M1b w/ bean dump   172
    2 circular beam dumps + spacer (17g each)   34
    Silicium beam dump w/o glass beam dump 490  
    Silicium beam dump w/ glass beam dump   501
    Rectangular small beam dump with graduated base (173 g each) x6   1038
    Rectangular small beam dump with fork (212 g each) x1   212
    Rectangular large beam dump with graduated base x1  

    195

    Total 15 196 12462

    Difference in weight= -2734g. We added weights of 935g, 943g, 480g, 221g and 223g and removed a small weight of 73g (2729g) in order to recover the same weight as before the intervention .

    We replaced B5_PD2 photodiode: Air Box 25 replaced by Air Box 56. B5_PD2 gains and current limit have been adjusted.

    We replaced B1p_PD2 photodiode: Air Box 1 replaced by Air Box 63.

    We installed a U-shape beam dump on the following photodiodes :

    • B5_PD1
    • B5_PD2
    • B1pP_PD1
    • B1p_PD1
    • B1p_PD2
    • B1_PD3
    • B1s_PD1
    • B1sP_PD1

    All photodiodes that have been removed from their base for the installation of the U-shape beam dump have been tested to check that Vbias can be put On and that the shutter is working.

    They have all been tested with light, all are working fine.

     

    Then we started with the beam dumps:

    We replaced B1p_M1b mount with the one with a beam dump glued on the side + a beam dump at the back.

     

    Finally we decided to change the splitters.

    • B5_M1 ==>  W216 (R=0.2) replaced by 36 (R=0.3)
    • B5_M4 ==> W208 (R=0.5) replaced by w228 (Rs=1, Tp=1). W208 had wedge on top, W228 now on the right
    • B5_M2 ==> W217 (R=0.2) replaced by Adv211 (R=0.5). W217 had wedge on top, Adv211 now on the right
    • B1p_M1==> W215 (R=0.2) replaced by Adv212 (R=0.5)
    • B1p_M4 ==> W204 (R=0.5) replaced by W215 (R=0.2). wedge changed from right to left
    • B1p_M2 ==> W203 (R= 0.5) relpaced by Adv209 (R=0.85)
    • B1_M1 ==> AdV210 (R=0.99) replaced by W209 (R=0.999)
    • B1p wedges have been reversed: B1p_PBS now on the left and B1p_M5 now on the right

    Circular beam dumps have been added in transmission of  B1s1_M2 and B1s1_M3 (inside the mount).

    We installed 8 glass beam dumps on the reflexion of lenses and ghost beams.

    We have displaced B1s_L2 towards east by approximately 20 mm.

    Alignmennt check are on-going.

     

      Comments to this report:
      bonnand, tournefier, pacaud. - 15:58 Thursday 10 June 2021 (52109) Print this report

      During the intervention on SDB2, we changed the malfunctionning netcom ethernet bridge.

      We followed the procedure in attachments.

      Unfortunately it was not pinging on the network, we asked Stefano Cortese to check if it was declared correctly on the network. It was the case but Stefano realized that the bridge was not in DHCP mode.

      After a few trials, we put the jumper at the back in the factory settings and press the "reset" button.

      Then it was pinging and we put back the jumper in the condition described in the document and we resetted the netcom bridge.

      Then we checked that we could control the quadrant photodiodes, it was the case.

      We also check that the netcom bridge is setting it up correctly when the ethernet switch is switch off and on (they share the same power supply).

      Everything work correctly now for the control of the quadrant photodiodes.

      Non-image files attached to this comment
      Search Help
      ×

      Warning

      ×